Which of the following statements is/are true? Explain why the false one is false. A. If u = (u1; u2) with ui not eqaual to 0 for some 1 <= i <= 2; then the unit vector in the direction of u is 1/sqrt( u1^2 +u2^2 )(u1; u2) : (2) B. u = (1; 2; 3) is the unit vector in the direction of u: (
1
Expert's answer
2020-05-04T19:52:05-0400

A) By the definition, unit vector: u' = u / ||u|| = (u1, u2) / (u12 + u22)1/2. True.

B) u' = u / ||u|| = (u1, u2, u3) / (u12 + u22 + u32)1/2 = (1, 2, 3) / (12 + 22 + 32)1/2 = (1/(14)1/2)(1, 2, 3).